Telephone calls are being placed through a certain exchange at random times on the

average of four per minute. Assuming a Poisson process, determine the probability

that in a 15-second intervals, there are 3 or more calls.


1
Expert's answer
2021-06-01T08:59:20-0400

"P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

λ=\frac{4}{4}=1 \\

P(X≥3) = 1- P(X\leq 2) = 1-[P(X=0)+P(X=1)+P(X=2)] \\

= 1- [\frac{e^{-1}(1)^0}{0!}+\frac{e^{-1}(1)^1}{1!}+ \frac{e^{-1}(1)^2}{2!}] \\

= 1-e^{-1}(1+1+\frac{1}{2}) \\

=1-2.5 \times e^{-1} \\

=1-0.9175 \\

=0.0825"